OpenStudy (anonymous):

what expression is equivalent to x2/7

OpenStudy (jdoe0001):

\(\bf \large { a^{\frac{{\color{blue} n}}{{\color{red} m}}} \implies \sqrt[{\color{red} m}]{a^{\color{blue} n}} \qquad \qquad \sqrt[{\color{red} m}]{a^{\color{blue} n}}\implies a^{\frac{{\color{blue} n}}{{\color{red} m}}}\qquad thus \\ \quad \\ x^{\frac{{\color{blue}{ 2}}}{{\color{red}{ 7}}}}\implies ? }\)